### v1

@@ 5.1 Evaluation metrics
${{\rm ACC}(v,l)=\frac{{\textstyle\sum_{m,n}}L(m)\hat{X}_{pred}(l)[v,m,n]\hat{X}_{true}(l)[v,m,n]}{\sqrt{{\textstyle\sum_{m,n}}L(m)({\hat{X}_{pred}(l)[v,m,n]})^{2}{\textstyle\sum_{m,n}L(m)({\hat{X}_{true}(l)[v,m,n]})^{2}}}}},$ (9)
where ${\hat{X}_{pred/true}(l)[v,m,n]}$ represents the long-term-mean-subtracted value of predicted or true variable $v$ at the location denoted by the grid co-ordinates $(m,n)$ at the forecast time-step $l$. The long-term mean of a variable is just the mean value of it over a large number of historical samples in the training dataset. The long-term mean-subtracted variables ${\hat{X}_{pred/true}}$ represent the anomalies of those variables that are not captured by the long term values. $L(m)$ is the latitude weighting factor at the co-ordinate $m$ which is defined in Eq. (8). The latitude-weighted RMSE for a forecast variable $v$ at forecast time-step $l$ is defined by the following equation:
${{\rm RMSE}(v,l)=\sqrt{\frac{1}{NM}\sum_{m=1}^{M}\sum_{n=1}^{N}L(m){(X_{pred}(l)[v,m,n]-X_{true}(l)[v,m,n])}^{2}}},$ (10)
where $X_{pred/true}(l)[v,m,n]$ represents the value of predicted or true variable $v$ at the location denoted by the grid co-ordinates $(m,n)$ at the forecast time-step $l$.
@@ 5.2 Comparison with state-of-the-art models
We compare the forecast performance of EWMoE with FourCastNet, ClimaX and Pangu-Weather, three leading deep learning methods for global weather forecasting. Figures 2 and Figure 3 evaluate different methods on forecasting four key weather variables at lead time from 1 to 8 days in terms of ACC and RMSE, respectively. The results show that EWMoE has both higher ACC and lower RMSE than FourCastNet [12] and ClimaX [13] for all the variables analyzed. For short-range forecasting (1-3 days), EWMoE demonstrates a comparable level of forecasting accuracy as Pangu-Weather [14]. In addition, as the forecast time increases, significant improvement with our EWMoE is observed and EWMoE outperforms Pangu-Weather from day 3, demonstrating EWMoE’s remarkable ability and stability for short-to-medium-range weather forecasting.
Moreover, we also note that EWMoE achieves this strong performance with much less training data and computing resources compared with the baselines. We train our EWMoE on 2 years of training data, which is approximately 18$\times$ less data than FourCastNet and ClimaX’s 37 years of training data, and 120$\times$ less than that used for Pangu-Weather, which uses 39 years of training data with 13 pressure levels. The training of EWMoE was completed under 9 days on 2 Nvidia 3090 GPUs. In contrast, FourCastNet took 16 hours to train on 64 A100 GPUs, ClimaX took 7 days on 80 V100 GPUs, and Pangu-Weather required 64 days on 192 V100 GPUs. Our novel weather forecasting MoE model generates accurate forecasts with much less training data and computational cost, which will facilitate future works that build upon our proposed framework.
@@ 5.4 Ablation studies
Effect of meteorology-specific embedding. We conduct experiments to compare the performance of EWMoE with and without meteorology-specific embedding to evaluate the its effectiveness in extracting the internal relations between meteorology variables. Figure 6(a) shows the superior performance of meteorology-specific embedding compared with the standard ViT embedding at all forecast time, indicating that it is a crucial component in capturing interactions between atmospheric variables through cross-attention.
(a) Effect of meteorology embedding.
@@
Effect of MoE layer. We evaluate the effectiveness of the MoE layer in EWMoE. As shown in Figure 6(b), EWMoE with the MoE layer significantly outperforms model with a standard feed-forward network, and the performance gap becomes larger as the forecast time increases. We attribute this result to the ability of the MoE layer, which allows increasing the model capacity and flexibility. We also note that EWMoE achieves this improvement without increasing computing requirements, as only a small portion of parameters are activated during training. This suggests that applying MoE structure to weather forecasting is promising. Moreover, we also conduct extensive experiments to evaluate the importance of auxiliary loss used in the MoE layer routing and the number of selective top-$k$ experts, as shown in Figure 7.
@@ 6 Conclusion
In this paper, we introduce EWMoE, an advanced and effective deep learning model for weather forecasting. By integrating three novel components, meteorology-specific embedding, an MoE layer and two specific loss functions, it excels at a resolution of 0.25 and forecast time of up to 8 days, outperforming the leading models such as FourCastNet and ClimaX, and competing well with Pangu-Weather in short-range forecasting. It is worth mentioning that EWMoE achieves this superior performance with significantly less training data and computing resources, addressing the challenges of computational efficiency and prediction accuracy. Our study also provides insights for modeling the interactions among atmospheric variables, demonstrating the feasibility and potential of implementing the MoE paradigm in weather forecasting tasks. We hope that our work will inspire future work on applying effective MoE architecture to a wider range of climate researches.

### v2

@@ 5.2 Evaluation metrics
${{\rm ACC}(v,l)=\frac{{\textstyle\sum_{m}}L(m)\hat{X}_{pred}\hat{X}_{true}}{\sqrt{{\textstyle\sum_{m}}L(m)({\hat{X}_{pred}})^{2}{\textstyle\sum_{m}L(m)({\hat{X}_{true}})^{2}}}}},$ (9)
where ${\hat{X}_{pred/true}}$ represents the long-term-mean-subtracted value of predicted or true variable $v$ at the location denoted by the grid co-ordinates at the forecast time-step $l$. The long-term mean of a variable is just the mean value of it over a large number of historical samples in the training dataset. The long-term mean-subtracted variables ${\hat{X}_{pred/true}}$ represent the anomalies of those variables that are not captured by the long term values. $L(m)$ is the latitude weighting factor at the co-ordinate $m$ which is defined in Eq. (8). The latitude-weighted RMSE for a forecast variable $v$ at forecast time-step $l$ is defined by the following equation:
${{\rm RMSE}(v,l)=\sqrt{\frac{1}{NM}\sum_{m}^{M}\sum_{n}^{N}L(m){(X_{pred}-X_{true})}^{2}}},$ (10)
where $X_{pred/true}$ represents the value of predicted or true variable $v$ at the location denoted by the grid co-ordinates at the forecast time-step $l$.
@@ 5.3 Comparison with state-of-the-art models
We compare the forecast performance of EWMoE with FourCastNet, ClimaX, Pangu-Weather and GraphCast, four leading deep learning methods for global weather forecasting. Figures 2 and Figure 3 evaluate different methods on forecasting four key weather variables at lead time from 1 to 8 days in terms of ACC and RMSE, respectively. The results show that EWMoE has both higher ACC and lower RMSE than FourCastNet [12] and ClimaX [13] for all the variables analyzed. For short-range forecasting (1-3 days), EWMoE demonstrates a comparable level of forecasting accuracy as Pangu-Weather [14]. In addition, as the forecast time increases, significant improvement with our EWMoE is observed and EWMoE outperforms Pangu-Weather from day 3, demonstrating EWMoE’s remarkable ability and stability for short-to-medium-range weather forecasting. Compared with GraphCast, each model has its own advantages. In terms of the ACC metric, EWMoE performs better than GraphCast, while in terms of the RMSE metric, GraphCast is slightly better than EWMoE. This result may be attributed to the fact that GraphCast uses a 12-step autoregressive finetuning strategy to reduce the error accumulation in long lead predictions but increases the consumption of training resources at the same time.
Moreover, we note that EWMoE achieves this strong performance with much less training data and computing resources compared with the baselines. We train our EWMoE on 2 years of training data, which is approximately 18$\times$ less data than FourCastNet and ClimaX’s 37 years of training data, and 120$\times$ less than that used for Pangu-Weather and GraphCast, which use 39 years of training data with 13 pressure levels. The training of EWMoE was completed under 9 days on 2 Nvidia 3090 GPUs. In contrast, FourCastNet took 16 hours to train on 64 A100 GPUs, ClimaX took 7 days on 80 V100 GPUs, Pangu-Weather took 64 days on 192 V100 GPUs and GraphCast took 4 weeks on 32 Google Cloud TPU v4 devices. Our novel weather forecasting MoE model generates accurate forecasts with much less training data and computational cost, which will facilitate future works that build upon our proposed framework.
@@ 5.5 Ablation studies
Effect of 3D absolute position embedding. We conduct experiments to compare the performance of EWMoE with and without 3D absolute position embedding to evaluate its effectiveness in extracting the 3D geographical location features. Figure 6(a) shows the superior performance of 3D absolute position embedding compared with the standard ViT position embedding at all forecast time, indicating that it is a crucial component in modeling geographical characteristics of different meteorological variables.
(a) Effect of 3D absolute position embedding.
@@
Effect of the MoE layer. We evaluate the effectiveness of the MoE layer in EWMoE. As shown in Figure 6(b), EWMoE with the MoE layer significantly outperforms model with a standard feed-forward network, and the performance gap becomes larger as the forecast time increases. We attribute this result to the ability of the MoE layer, which allows increasing the model capacity and flexibility. The total number of model parameters with only one FFN layer is 43 million. After using the MoE layer with 20 experts, the total number of model parameters has reached to 580 million. The total number of model parameters has increased by 13 times, allowing EWMoE to better extract and model meteorological data features. We also note that EWMoE achieves this improvement without increasing computing requirements, as only a small portion of parameters are activated during training. This suggests that applying MoE structure to weather forecasting is promising. Moreover, we also conduct extensive experiments to evaluate the importance of auxiliary loss used in the MoE layer routing and the number of selective top-$k$ experts, as shown in Figure 7.
@@ 6 Conclusion
In this paper, we introduce EWMoE, an advanced and effective deep learning model for weather forecasting. By integrating three novel components, 3D absolute position embedding, an MoE layer and two specific loss functions, it excels at a resolution of 0.25 and forecast time of up to 8 days, outperforming the leading models such as FourCastNet and ClimaX, and competing well with Pangu-Weather and GraphCast in short-range forecasting. It is worth mentioning that EWMoE achieves this superior performance with significantly less training data and computing resources, addressing the challenges of computational efficiency and prediction accuracy. Our study also provides insights for modeling the interactions among atmospheric variables, demonstrating the feasibility and potential of implementing the MoE paradigm in weather forecasting tasks. We hope that our work will inspire future work on applying effective MoE architecture to a wider range of climate researches.